OpenStudy (anonymous):

convert x^2 -x -8 into vertex form

hartnn (hartnn):

you know how to complete the square ?

OpenStudy (anonymous):

factor?

hartnn (hartnn):

whatst he co-efficient of x ?

hartnn (hartnn):

for \(x^2+bx+c\) the co-efficient of 'x' is 'b' so to complete the square, we add and subtract, \((1/2)b^2\) s, what do we add and subtract here ?

OpenStudy (anonymous):

(1/2)1^2?

OpenStudy (anonymous):

i mean (1/2)(-1)^2

OpenStudy (anonymous):

which is just .5

hartnn (hartnn):

sorry, i meant \((1/2\times b)^2 = (-1/2)^2 =1/4\) got this ? add and subtract this!

hartnn (hartnn):

\((x^2-x+1/4 ) + (-8-1/4)\)

hartnn (hartnn):

now first 3 terms form a perfect square, can you figure out ? \((..-..)^2\)

hartnn (hartnn):

\((x- ...?...)^2 \)

OpenStudy (anonymous):

ohh

OpenStudy (anonymous):

umm

hartnn (hartnn):

1/4 is the square of which number ?

OpenStudy (anonymous):

1/2

OpenStudy (anonymous):

(x-1/2)(x-1/2) ?

hartnn (hartnn):

yes correct! so we have \(\Large (x-1/2)^2 +(-1/4-8)\) now just simplify (-1/4-8)

OpenStudy (anonymous):

-8.25

hartnn (hartnn):

find in terms of fraction

OpenStudy (anonymous):

-33/4?

hartnn (hartnn):

thats right! so your vertex form is just : \(\Large (x-1/2)^2 -33/4\) thats it! :)

OpenStudy (anonymous):

omg! wow thanks so much

hartnn (hartnn):

welcome ^_^